OpenStudy (anonymous):

What is the volume of this composite solid? A)588in3 B)784in3 C)915in3 D)980in3

OpenStudy (anonymous):

Plz help I need to finish my homework

OpenStudy (littlebird):

You need to give more information about the solid.

OpenStudy (anonymous):

ok hold on.

OpenStudy (anonymous):

here you go Littlebird

OpenStudy (anonymous):

are you still workin it out.Im not trying to rush u or anything just asking

OpenStudy (littlebird):

The shape is made of three different 3D shapes. The cubes would use the formula Volume= basexwidthxheight. For the triangle you would find the surface area of one of its triangular sides and multiply that by its height.

OpenStudy (mathstudent55):

This is easier than you think. There are two prisms that are 8 in. x 7 in. x 7 in. And one piece that is half of a prism 8 in. x 7 in. x 7 in.

OpenStudy (anonymous):

ok thx. can you walk me through it

OpenStudy (anonymous):

ohh never mind thank you littlebird and mathstudent55:)

OpenStudy (littlebird):

Ok :)

OpenStudy (anonymous):

Can you guys help me with another one plz. here it is: Which of the formulas below would be used to calculate the volume of the composite figure below? A.V =1/2BH t h +/wh B. V =/wh +s3 C. V= pier 2h +1/2bh t h D. V = 1/3/wh + /wh

OpenStudy (mathstudent55):

V = LWH V = 8 in. * 7 in. * 7 in. + 8 in. * 7 in. * 7 in. + (8 in. * 7 in. * 7 in.)/2 V = 980 in^3

OpenStudy (anonymous):

thx can you guys help me with the other question tht I posted plz

OpenStudy (mathstudent55):

The upper part is the area of a triangle times a height. The lower part is a length times a width times a height.

OpenStudy (littlebird):

Whichever answer has both the formula for a rectangular prism and a triangle shape is the one.

OpenStudy (anonymous):

ok thank yoy so much for helping :) Have a nice afternoon

OpenStudy (mathstudent55):

You're welcome. Thanks, you too.

OpenStudy (anonymous):

which letter is it

OpenStudy (anonymous):

mathstudent55 which letter is it

OpenStudy (mathstudent55):

I don't understand your choices. What does pier mean? What is /wh?

OpenStudy (anonymous):

pier means the pie sign and r means radius of circle. lwh / length of prism,width of prism and hieght of prism

OpenStudy (mathstudent55):

Ok. Got it. There is no circle here, so the answer has no pi in it. Also there is a triangle here, so the answer must have (1/2)bh.

OpenStudy (anonymous):

ok thx.

OpenStudy (mathstudent55):

wlcm